Schedule of Cash Payments for a Service Company

EastGate Physical Therapy Inc. is planning its cash payments for operations for the first quarter (January–March). The Accrued Expenses Payable balance on January 1 is $34,300. The budgeted expenses for the next three months are as follows:

January February March
Salaries $78,900 $96,000 $106,300
Utilities 6,500 7,200 8,600
Other operating expenses 60,000 65,400 72,000
Total $145,400 $168,600 $186,900

Other operating expenses include $4,300 of monthly depreciation expense and $1,000 of monthly insurance expense that was prepaid for the year on May 1 of the previous year. Of the remaining expenses, 70% are paid in the month in which they are incurred, with the remainder paid in the following month. The Accrued Expenses Payable balance on January 1 relates to the expenses incurred in December.

Prepare a schedule of cash payments for operations for January, February, and March. Enter all amounts as positive numbers.

EastGate Physical Therapy Inc.
Schedule of Cash Payments for Operations
For the Three Months Ending March 31
January February March
$ $ $
Total cash payments $ $ $
Answer #1
EastGate Physical Therapy Inc.
Schedule of Cash Payments for Operations
For the Three Months Ending March 31
January February March
Payments of prior month's expense 34300 42030 48990
Payments of current month's expense 98070 114310 127120
Total cash payments 132370 156340 176110
Formulas used:
Payments of prior month's expense 34300(given) =(145400-4300-1000)*30% =(168600-4300-1000)*30%
Payments of current month's expense =(145400-4300-1000)*70% =(168600-4300-1000)*70% =(186900-4300-1000)*70%